the electric potential v at any point x,y,z in space is given by v=4x^2 volt then the electric field at the point 1,0,2 v/m is

Grade:12

1 Answers

Riddhish Bhalodia
8 years ago
E = (-\frac{\partial V}{\partial x},-\frac{\partial V}{\partial y}, -\frac{\partial V}{\partial z})
so here we have
E = (-8x,0,0)
so at 1,0,2
E = (-8,0,0)